OpenStudy (anonymous):

At the end of the first day, 7 weeds appear in your neighborhood park. Each day, the number of weeds increases by four times. How many weeds will be in the park at the end of 14 days? please helpp

OpenStudy (anonymous):

i dont know what formula to use

OpenStudy (anonymous):

i just like did the work and wrote it all out, like (day 1=7), (day2= 7*4), (day3=7*4*4).... but i got 469762048... and thats not right

OpenStudy (anonymous):

35,456,231 626,349,395 71 392 these are the choices

OpenStudy (radar):

first day 7 2nd day 7*4 3rd dat 7*4*4 4th day 7*4*4*4 14 day 7 *4^13=469,762,048 I got that too. hopefully someone will show me where I veered of the path lol.

OpenStudy (anonymous):

haha ya i did it twice and im almost positive the right answer isnt a choice

OpenStudy (anonymous):

i just guessed 626,349,395, cuz its the closest lol

OpenStudy (anonymous):

i found the answer if your curious.. http://answers.yahoo.com/question/index?qid=20120302205322AAOZKd3 its on that page

OpenStudy (radar):

I was wondering, there is 7 weeds on day one, would the increase of 4*7 be added to the previous day number like 7+28=35 .......??????? I would go for the biggest number there as in my opinion that is the best answer

OpenStudy (radar):

O.K. I looked it up, I guess I had veered off the path.

OpenStudy (anonymous):

thanks anyways though